OpenStudy (anonymous):

Which of these sets of numbers contains no irrational numbers? (1 point) , , –8.15 ,

OpenStudy (anonymous):

what sets?

OpenStudy (anonymous):

square root -5 ,- square root 196 , –8.15 or square root -144, square root 49, 6.6 or -2. 3/8 square root 10, -0.6868... or -1, 5/6 square root 11 ,

OpenStudy (anonymous):

im really sorry, but thats still not very clear to me, are those negative signs? i want to help but i dont understand what the sets are

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the first "set" is \[\sqrt{-5}\] ?

OpenStudy (anonymous):

that and -196 and -8.15 then the others are separated by the word "or"

OpenStudy (anonymous):

aah i see, ok i'll get on it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

are you sure its \[\sqrt{-5} \text{ instead of } -\sqrt{5}\] because the first is a "complex" number, the square root of a negative i'll continue anyway: square root -5 ,- square root 196 , –8.15 : \[\sqrt{-5} \text{ , } -\sqrt{196} \text{ , } -8.15 \] a number is irrational if it cannot be written as a fraction although the \[- \sqrt{196} = - \sqrt{14^2} = -14\] which is rational, \[\sqrt{-5} = i \sqrt{5}\] which cannot be written as a fraction as the square root of 5 is a surd

OpenStudy (anonymous):

are you sure there are all those negative signs? it seems a bit out of context compared with the rest of the question

OpenStudy (anonymous):

\[-\sqrt{5}\]

OpenStudy (anonymous):

that makes much more sense

OpenStudy (anonymous):

sorry helping my son with this and I have never did this kind of math

OpenStudy (anonymous):

no problem : ) i'll just write it out with the equation thingy, then i'll explain -> square root -5 ,- square root 196 , –8.15 or square root -144, square root 49, 6.6 or -2. 3/8 square root 10, -0.6868... or -1, 5/6 square root 11 : \[[ -\sqrt{5} , -\sqrt{196} , -8.15 ] \text{ , } [ -\sqrt{144}, \sqrt{49} , 6.6] , [-2,\frac{3}{8}\sqrt{10},-0.6868...] , [-1,\frac{5}{6} \sqrt{11} ]\]

OpenStudy (anonymous):

ah a bit got cut off

OpenStudy (anonymous):

i'll go ahead and do some explaining in maths we have "rational numbers" which is a facy way of saying "numbers that can be written as fractions" "irrational numbers" are numbers that cant be written as fractions any counting number, positive or negative, can be written as a fraction eg \[1 = \frac{1}{1} \text{ , } 15 = \frac{15}{3} \text{ , } -6 = \frac{-6}{1}\] other examples of rationals: \[ \frac{3}{8} \text{ , } \frac{4}{10} \text{ , } \frac{22}{27}\]

OpenStudy (anonymous):

some numbers we cannot write as fractions, for example \[\sqrt{2} \] cannot be written as a fraction

OpenStudy (anonymous):

oooooo ic thank you

OpenStudy (anonymous):

here is a good video that may help: http://www.youtube.com/watch?v=QIoVPtbEUjw

OpenStudy (anonymous):

if we take a square root of a square number, that is rational eg sqrt(4) = 2 = 2/1

OpenStudy (anonymous):

sorry i made a mistake earlier, 15 = 15/1 not 15/3